OpenStudy (anonymous):

10. The first of two movies lasted 8 minutes less than three times as long as the second. Together the two movies lasted 156 minutes. How long was the second movie? A. 39 minutes C. 43 minutes B. 41 minutes D. 52 minutes

OpenStudy (hoblos):

x= 3y-8 x+y = 158 3y-8 +y = 158 4y = 166 y=41.5

OpenStudy (hoblos):

it is 41.5 min so the best answer is B

OpenStudy (anonymous):

So the answer is 41?

OpenStudy (anonymous):

i can't see the answer @Luis Rivera what did you put?

OpenStudy (hoblos):

im sorryyy it is exactly 41 so it is B i miss read it first

OpenStudy (anonymous):

It won't let me. what u put?

OpenStudy (hoblos):

x= 3y-8 x+y = 156 3y-8 +y = 156 4y = 164 y=41